A sample of n = 16 observations is drawn from a
normal population with μ = 1,000 and σ = 200.
Find the following.
a. P(X > 1,050)
b. P(X < 960)
c. P(X > 1,100)
1
Expert's answer
2020-12-08T09:02:09-0500

Solution

a. P(x>1050)


"P(X>1050) =1-P(X<1050)"

"=1- \phi ({x- \mu \over {\sigma \over \sqrt{n}}})"

"=1 - \phi ({1050-1000 \over {200 \over \sqrt{16}} })=1- \phi (1)""=1-0.8413""=0.1587"

b. P(X<960)


"=\phi ({960-1000 \over {200 \over \sqrt{16}}} )= \phi(-0.8)"

"=0. 2119"

c. P(X>1100)


"=1 - \phi ({1100-1000 \over {200 \over \sqrt{16}} })=1- \phi (2)"

"=1-0.9772""=0.0228"